# DiffDA: a diffusion model for weather-scale data assimilation

### 11 Jan 2024 v1

@@ 2.5 Selection of diffusion model structure
Our method provides a lot of flexibility on the selection of the neural network structure for $\epsilon_{\theta}(\mathbf{x}^{j}_{i},\hat{\mathbf{x}}_{i},j):\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{N}^{+}\rightarrow\mathbb{R}^{n}$ as any neural network that matches the function signature will work. However, $n$ can be a large number of tens of millions in practice rendering many neural network architectures infeasible due to resource constraints. It is crucial for a neural network to utilize the spatial information in the state vector $\mathbf{x}$ in order to learn efficiently.
Instead of creating a new architecture, we employ a proven architecture in ML weather forecast models which has a similar signature $\mathbb{R}^{c\cdot n}\rightarrow\mathbb{R}^{n},c\in\mathbb{N}^{+}$. Due to this similarity, neural networks that perform well in forecasting should also do well in data assimilation task, and it is likely to take less training steps using pretrained weights of the forecast model when training the diffusion model. Moreover, thanks to this, we can adapt the neural network structure with ease from any state-of-the-art weather forecast model.
@@ 3.1 Implementation
We demonstrate our method in a real-world scenario with a horizontal resolution of 0.25 degree and 13 vertical levels. This matches the resolution of the WeatherBench2 dataset [Rasp et al., 2023] used by state-of-the-art ML weather forecast models. We take the GraphCast model as the backbone of the diffusion model because the pretrained model takes in states at two consecutive time steps $\mathbf{x}_{i-1},\mathbf{x}_{i}$ to predict $\mathbf{x}_{i+1}$ which takes much less effort than other forecast models to re-purpose it to $\epsilon_{\theta}(\mathbf{x}^{j}_{i},\hat{\mathbf{x}}_{i},j)$ given that there is one-to-one matching between $\mathbf{x}_{i-1}$ and $\mathbf{x}^{j}_{i}$, as well as $\mathbf{x}_{i+1}$ and $\hat{\mathbf{x}}_{i}$. As is determined by the pretrained GraphCast model, the input size $n$ is set to $(6\times 13+5)\times 721\times 1440$ where there are 6 pressure level variables, 5 surface level variables, 13 pressure levels, 721 latitude and 1440 longitude grid points on a longitude-latitude horizontal grid.
The diffusion model is implemented with the jax library [Bradbury et al., 2018] and the official implementation of GraphCast. We use the AdamW optimizer [Loshchilov and Hutter, 2018] with a warm-up cosine annealing learning rate schedule and data parallel training on 80 AMD MI250 GPUs with a (global) batch size of 80 for 12 epochs. Gradient checkpoints are added in the GraphCast model to reduce the GPU memory footprint.
@@ 3.2 Training data
We use the WeatherBench2 dataset as the first part of training data. The dataset contains values for our target atmospheric variables from 1979 to 2016 with a time interval of 6 hours extracted from the ERA5 reanalysis dataset. Data ranging from 1979 to 2015 is used for training while the rest is used for validation and testing. The second part of the training data is generated by running 48-hour GraphCast forecasts (with 8 time steps) using the data from the first part as initial conditions. Then, the two parts are paired up according to their physical time. Before feeding it to the model, the input data is normalized using the vertical-level-wise means and standard deviations provided by the pretrained GraphCast model. The output of the model is de-normalized in a similar way. Due to the fact that the predicted state $\hat{\mathbf{x}}$ is largely close to the ground truth $\mathbf{x}^{*}$, we add a skip connection from $\hat{\mathbf{x}}$ to the de-normalized model output and use the normalized difference $\mathbf{x}^{0*}_{i}-\hat{\mathbf{x}},\hat{\mathbf{x}}$ as the actual input data to the diffusion model. Following GraphCast, the normalization and de-normalization use different sets of means and standard-deviations for original data $\hat{\mathbf{x}}$ and delta data $\mathbf{x}^{0*}_{i}-\hat{\mathbf{x}}$.
@@ 3.3 Treatment of conditioning for sparse observations
Acknowledging the multidimensional nature of the state vector and that most meteorological observations are co-located horizontally (longitude and latitude), we opt for a simplified setting in the conditioning of sparse observations. In this scenario, the observational data is $m$ sampled columns of the ground truth state vector with $6\times 13+5$ values in each column: $\mathbf{y}\in\mathbb{R}^{(6\times 13+5)\times m}$. The mask is simplified to a 2D mask $\mathbf{m}_{s}\in\mathbb{R}^{721\times 1440}$ which is broadcast to other dimensions when doing point-wise multiplication with the state vector in Equation (missing) 4. Interpolation is also applied independently over 2D horizontal slices for each variable and level. To interpolate appropriately for values on a sphere, we project the 2D slice into 3D space on a unit sphere and use the unstructured interpolation function griddata from the scipy package [Virtanen et al., 2020].
@@ 3.4 Experiment settings and results
Figure 3: Overview of the experiment settings. Hexagon represents atmosphere states, black arrow represents data assimilation, brown solid arrow represents 6h forecast, brown dashed arrow represents 48h forecast, hexagon with dashed edges and sparse points represent sparse observations, and wide arrow represents target and reference in each experiment. Forecast using GraphCast takes in atmosphere states at two time steps, this is ignored in this diagram for simplicity.
@@
We demonstrate the effectiveness of our method by performing real-world inspired experiments with increasing complexity (Figure 3). In the basic scenario, we perform data assimilation on 48h forecast and observation data, then directly compare the assimilated data with the ground truth data. Furthermore, we evaluate our method with autoregressive data assimilation where the predicted state is produced by a 6h forecast from the assimilated state 6 hour ago. We design this experiment to test whether the assimilate data will deviate from the ground truth data, which is crucial in real world applications. We also designed another experiment to compare the 48h forecast result based on assimilated data and ground truth data, and evaluate the effect of data assimilation method on the forecast skill.
In all the experiments, observations are simulated by taking random columns from the ERA5 dataset considered as the ground truth. We vary the number of observed columns $m$ in the experiments from 1,000 to 40,000 to test the convergence property of our data assimilation method.
@@
Figure 4: Root mean square error (RMSE) of geopotential at 500hPa, temperature at 850hPa, and temperature at 2m from the single step assimilated data and GraphCast forecast data (shown by the numbers in the cell). The error is calculated against the ERA5 data. The cells are color-coded with the RMSE relative to the 6h forecast RMSE where green means better and red means worse.
Single step data assimilation
@@
In this experiment, we perform data assimilation with a 48h GraphCast forecast and observed columns from 1,000 to 40,000, calculate the error between the assimilated data and the ground truth data, and compare this error against forecast errors with lead times from 6h to 48h.
@@
The result is presented in Figure 4 where we pick three representative variables closely related with forecast skills [Ashkboos et al., 2022, Rasp et al., 2020] including geopotential at 500hPa (z500), temperature at 850hPa (t850), and temperature at 2m (t2m) as the variable. With only 1,000 observed columns ($<0.1\%$ total columns), the assimilated data already achieves a lower RMSE compared to the input 48h forecast data for z500 and t850, while the error of t2m is even lower than the 6h forecast. The error further decreases as the number of observed columns increases. On the other side of the spectra, with 40,000 observed columns ($<3.9\%$ total columns), the RMSE of all three variables are lower than the 6h forecast error. This indicates our assimilation method can indeed improve over the given predicted state and converges to the ground truth as the number of observed columns increases.
@@
In an operational data assimilation system, data assimilation is performed in an autoregressive way: assimilated data from previous time steps are used for producing forecast at the next time step which is in turn used for assimilating data at that time step. Following the operational setting, we perform autoregressive data assimilation cycle: it starts from a 48h forecast at 0h, and repeat the data assimilation - 6h forecast cycle autoregressively. In the data assimilation cycle, two strategies of sampling the columns are used. One is to sample the columns at the same locations (Figure 5). The other resamples data at different locations in each iteration (Figure 6). The strategies helps estimate the performance in real-world cases since the reality is the mix of the two strategies where some observations are measured at the same locations such as weather stations and geostationary satellites, while locations of other observations change with time. We evaluate the autoregressive data assimilation result similar to the previous experiment.
In the results, we observed distinct patterns for fixed and non-fixed sampling strategies. With fixed sampling (Figure 5), the initial five iterations showed promising performance, but an error began to accumulate in subsequent iterations. Notably, this error demonstrated a diurnal pattern in the assimilation error for temperature at 2 meters (t2m). In contrast, the non-fixed sampling strategy showed a slower rate of error accumulation compared to fixed sampling (Figure 6). However, a significant issue was identified in both strategies: the assimilated error (solid lines) consistently exceeded the interpolation error (dash lines) after several iterations. This outcome suggests that our assimilation method is not effectively utilizing the input observational data to its full potential.
@@
Figure 5: RMSE of geopotential at 500hPa, temperature at 850hPa and temperature at 2m from autoregressively assimilated data (solid lines) and interpolated observations (dash lines). The observations at made at the same locations in each assimilation step.
@@
Figure 6: RMSE of geopotential at 500hPa, temperature at 850hPa and temperature at 2m from autoregressively assimilated data (solid lines) and interpolated observations (dash lines). The observations are made at different locations in each assimilation step.
Forecast on single step data assimilation result
@@
Since assimilated data is often used as the input of weather forecasts, it is important to test the forecast error using our assimilated data. In this experiment, we perform data assimilation at 0h, perform a 48h GraphCast forecast, then compare the forecast error again error of forecasts with varying lead times using ERA5 at 0h as inputs.
@@
Figure 7: Root mean square error (RMSE) of geopotential at 500hPa, temperature at 850hPa, and temperature at 2m from the 48h forecast using assimilated data as inputs and forecasts with lead times from 48h to 78h using ERA5 as inputs (shown by the numbers in the cell). The error is calculated against the ERA5 data. The cells are color-coded with the RMSE relative to the 72h forecast RMSE where green means better and red means worse.
@@
As the score board plot in Figure 7 shows, results from this experiment reveal that the 48h forecast error using assimilated data inputs gradually converges to the 48h forecast errors using ERA5 inputs. The error using assimilated data cannot be lower than using ERA5 inputs because the ERA5 data is used to emulate observations in the experiment. However, this setup allows us to compare the forecast error against forecasts with longer lead times to determine the extent of lead time lost due to the use of assimilated data as forecast inputs. With 8,000 observed columns, the error for z500, t850, and t2m is lower than the 72h forecast error. This implies that the lead time lost is less than 24 hours. With 40,000 observed columns, the error for z500 and t850 is approximately equivalent to the 54h forecast error. However, the t2m error remains larger than the 60h forecast error, potentially due to diurnal effects that cause a decrease in the forecast error of t2m from a 48h to a 60h lead time.
@@ 3.5 Conclusion
In summary, we propose a data assimilation method capable of assimilating high-resolution atmospheric variables using predicted state and sparse observations. Our approach is based on the denoising diffusion model, which is able to sample from conditional probability distributions. Notably, we adapted the pretrained GraphCast weather forecast model as the denoising model motivated by its compatible input and output shapes. Our method’s flexibility also allows for the possibility of integrating other forecast models, ensuring ease of updates and maintenance.
A key feature of our method is the enforcement of conditioning on the predicted state during both training and inference periods, and the conditioning of sparse observations exclusively at the inference stage. An additional benefit of this conditioning approach is the automatic generation of a post-processing model, should observations not be supplied at inference time.
Our experimental results validated the effectiveness of this approach: we observed that the assimilated data converges to the observations. When used as an input for forecast models, assimilated data resulted in a maximum lead time loss of 24 hours. Furthermore, when provided with sufficient observations, our method proved to be capable of producing reanalysis data.
Limitations
Currently, our method lacks a 4D assimilation capability and is unable to assimilate observations other than direct point measurements of assimilated variables. We are working on assimilating point observations at multiple time steps as well as general observations such at satellite imagery and radar soundings. The columns of ERA5 that are used as mock observations within this study are easier to use as they hold less uncertainty when compared to real observations that suffer from measurement errors. However, there are overall many more observations available in real-world forecasts when compare to the number use in this paper.
4 Acknowledgements
This study received funding from the MAELSTROM project funded by the European High-Performance Computing Joint Undertaking (JU; grant agreement No 955513).

### 05 Mar 2024 v2

@@ 2.5 Selection of diffusion model structure
Our method provides a lot of flexibility on the selection of the neural network structure for $\epsilon_{\theta}(\mathbf{x}^{j}_{i},\hat{\mathbf{x}}_{i},j):\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{N}^{+}\rightarrow\mathbb{R}^{n}$ as any neural network that matches the function signature will work. However, $n$ can be tens of millions in practice, rendering many neural network architectures infeasible due to resource constraints. A neural network must utilize the spatial information in the state vector $\mathbf{x}$ to learn efficiently.
Instead of creating a new architecture, we employ a proven architecture in ML weather forecast models. It has a similar signature $\mathbb{R}^{c\cdot n}\rightarrow\mathbb{R}^{n},c\in\mathbb{N}^{+}$ as the diffusion model. Due to this similarity, neural networks that perform well in forecasting should also do well in the data assimilation tasks, and it is likely to take fewer training steps using pre-trained weights of the forecast model when training the diffusion model. Moreover, we can easily update the backbone of the diffusion model with the state-of-the-art weather forecast model.
@@ 3.1 Implementation
We demonstrate our method in a real-world scenario containing 6 pressure-level variables and 4 surface variables, with a horizontal resolution of ${}^{\circ}$ and 13 vertical levels. This matches the resolution of the WeatherBench2 dataset [Rasp et al., 2023] used by state-of-the-art ML weather forecast models. We use the GraphCast model as the backbone of the diffusion model because the pretrained model takes in states at two consecutive time steps $\mathbf{x}_{i-1},\mathbf{x}_{i}$ to predict $\mathbf{x}_{i+1}$. It takes much less effort than other forecast models to re-purpose it to $\epsilon_{\theta}(\mathbf{x}^{j}_{i},\hat{\mathbf{x}}_{i},j)$ given that there is one-to-one matching between $\mathbf{x}_{i-1}$ and $\mathbf{x}^{j}_{i}$, as well as $\mathbf{x}_{i+1}$ and $\hat{\mathbf{x}}_{i}$. As is determined by the pre-trained GraphCast model, the input size $n$ is set to $(6\times 13+4)\times 721\times 1440$.
The diffusion model is implemented with the jax library [Bradbury et al., 2018], diffusers library [von Platen et al., 2023], and the official implementation of GraphCast. We use the AdamW optimizer [Loshchilov and Hutter, 2018] with a warm-up cosine annealing learning rate schedule that starts from $10^{-5}$, peaks at $10^{-4}$ after 1/6 of total training steps, and ends at $3\times 10^{-6}$. We perform data-parallel training on 48 NVIDIA A100 GPUs with a (global) batch size of 48 for 20 epochs. Gradient checkpoints are added in the GraphCast model to reduce the GPU memory footprint. The training takes approximately 2 days. All the inference is performed on a single node with one A100 GPU, which produces an assimilated state in around 15 minutes.
@@ 3.2 Training data
We use the WeatherBench2 dataset as the first part of the training data. The dataset contains values for our target atmospheric variables from 1979 to 2016 with a time interval of 6 hours extracted from the ERA5 reanalysis dataset. Data from 1979 to 2015 is used for training, while the rest is used for validation and testing. The second part of the training data is generated by running 48-hour GraphCast predictions (with 8 time steps) using the data from the first part as initial conditions. Then, the two parts are paired up according to their physical time. Before feeding it to the model, the input data is normalized using the vertical-level-wise means and standard deviations provided by the pre-trained GraphCast model. The output of the model is de-normalized using the same means and standard deviations. Since the predicted state $\hat{\mathbf{x}}$ is close to the ground truth $\mathbf{x}^{*}$, we add a skip connection from $\hat{\mathbf{x}}$ to the de-normalized model output and use the normalized difference $\mathbf{x}^{0*}_{i}-\hat{\mathbf{x}},\hat{\mathbf{x}}$ as the actual input data to the diffusion model. Following GraphCast, the normalization and de-normalization use different sets of means and standard deviations for original data $\hat{\mathbf{x}}$ and delta data $\mathbf{x}^{0*}_{i}-\hat{\mathbf{x}}$.
@@ 3.3 Treatment of conditioning for sparse observations
Acknowledging the multidimensional nature of the state vector and that most meteorological observations are co-located horizontally (longitude and latitude), we opt for a simplified setting in the conditioning of sparse observations. In this scenario, the observational data is $m$ sampled columns of the ground truth state vector with $6\times 13+4$ values in each column: $\mathbf{y}\in\mathbb{R}^{(6\times 13+4)\times m}$. The mask is simplified to a 2D mask $\mathbf{m}_{s}\in\mathbb{R}^{721\times 1440}$ which is broadcast to other dimensions when doing point-wise multiplication with the state vector in Equation (missing) 4. Interpolation is also applied independently over 2D horizontal slices for each variable and level. To interpolate appropriately for values on a sphere, we project the 2D slice into 3D space on a unit sphere and use the unstructured interpolation function griddata from the scipy package [Virtanen et al., 2020].
@@ 3.4 Experiment settings and results
Figure 4: Overview of the experiment settings. Hexagons represent atmosphere states, black arrows represent data assimilation, brown solid arrows represent 6-hour prediction, brown dashed arrows represent 48-hour prediction, hexagons with dashed edges and sparse points represent sparse observations, and wide arrows point out targets and references to compare in each experiment.
@@
We demonstrate the effectiveness of our method by performing real-world inspired experiments with increasing complexity (Figure 4). In the basic scenario, we perform data assimilation on 48-hour forecast and observation data, then directly compare the assimilated data with the ground truth. Furthermore, we evaluate our method with autoregressive data assimilation, where the predicted state is produced by a 6-hour prediction from the assimilated state 6 hours ago. We designed this experiment to test whether the assimilated data will deviate from the ground truth data, which is crucial in real-world applications. We also designed another experiment to compare the 48-hour forecast based on assimilated data and ground truth data and evaluate the effect of the data assimilation method on the forecast skill. For single-step data assimilation and forecast on single-step data assimilation, we run 16 parallel experiments with different inputs and random seeds and present the result in average values. For the autoregressive data assimilation, we run each experiment once due to the limitations of computation resources. In all the experiments, observations are simulated by taking random columns from the ERA5 dataset considered as the ground truth. We vary the number of observed columns $m$ in the experiments to test the convergence property of our data assimilation method.
@@
Figure 5: Root mean square errors (RMSEs, shown by the numbers in the cell) of geopotential at 500hPa, temperature at 850hPa, and temperature at 2m from the single-step assimilated data, and from 6-hour to 48-hour GraphCast forecasts. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 48-hour forecast errors.
Single-step data assimilation
@@
In this experiment, we perform data assimilation from a 48-hour prediction using GraphCast and columns of observations ranging from 1,000 to 40,000, then calculate the error between the assimilated data and the ERA5 data. Finally, we compare this error against GraphCast forecast errors with 6-hour to 48-hour lead times.
@@
The result is presented in Figure 5 where we pick three representative variables closely related to forecast skills [Ashkboos et al., 2022, Rasp et al., 2020] including geopotential at 500hPa (z500), temperature at 850hPa (t850), and temperature at 2m (t2m) as the target variables. More results can be found in subsection A.2. With only 1,000 observed columns ($<0.1\%$ total columns), the assimilated data achieves lower RMSEs than the input 48-hour forecast for z500 and t850. At the same time, the errors of t2m are even lower than the 6-hour forecast. The errors further decrease as the number of observed columns increases. On the other side of the spectra, with 40,000 observed columns ($<3.9\%$ total columns), the RMSEs of all three variables are lower than the 6-hour forecast error. This indicates our assimilation method can improve over the given predicted state while remaining consistent with the observations.
@@
In an operational data assimilation system, data assimilation is performed in an autoregressive way: assimilated data from previous time steps are used for producing forecast at the next time step, which is in turn used for assimilating data at that time step. Following this setting, we perform an autoregressive data assimilation cycle: it starts from a 48-hour forecast at 0 hour, and repeats the data assimilation - 6-hour forecast cycle autoregressively. We train two separate diffusion models for this task, one dedicated to 48-hour forecast inputs and the other dedicated to 6-hour forecast inputs. In the data assimilation cycle, two strategies of sampling the simulated observation columns are used. One is to sample the columns at the same locations (Figure 6). The other resamples data at different locations in each iteration (Figure 7). The strategies help estimate the performance in real-world cases since the reality is the mix of the two strategies where some observations are measured at the same locations, such as weather stations and geostationary satellites, while locations of other observations change with time. We evaluate the autoregressive data assimilation result similar to the previous experiment.
We observe distinct patterns for fixed and non-fixed sampling strategies in the results. With fixed sampling (Figure 6), the initial five iterations show promising performance, but errors begin to accumulate in subsequent iterations. The assimilated data with 1,000 columns of observations ($<0.1\%$ total columns) quickly diverts from the ERA5 data especially for the z500 variable, but this is not the case with more observations. It implies that the given observations are not sufficient to balance the error accumulation in the forecast model. In contrast, the non-fixed sampling strategy shows a slower error accumulation rate than fixed sampling (Figure 7). However, an issue is identified in both strategies: the assimilated errors (solid lines) consistently exceed the interpolation errors (dashed lines) after several iterations. This is less of a concern in real-world scenarios where lots of the observations, such as satellite imagery and radar soundings, can not be explicitly interpolated into atmospheric variables. Notably, the assimilation errors of t2m indicate a diurnal pattern with a period of 24 hours. Since the interpolation errors remain stable over iterations, this diurnal pattern is likely introduced by the GraphCast forecast model.
@@
Figure 6: Root mean square errors (RMSEs) of geopotential at 500hPa, temperature at 850hPa and temperature at 2m from autoregressively assimilated data (solid lines) and interpolated observations (dashed lines). The observations are made at the same locations in each assimilation step.
@@
Figure 7: Root mean square errors (RMSEs) of geopotential at 500hPa, temperature at 850hPa and temperature at 2m from autoregressively assimilated data (solid lines) and interpolated observations (dashed lines). The observations are made at different locations in each assimilation step.
Forecast on single-step assimilated data
@@
Since assimilated data is often used as the input of weather forecasts, it is important to test the forecast error using our assimilated data. In this experiment, we perform data assimilation at 0 hour, perform a 48-hour prediction using GraphCast, and then compare the resulting errors against the forecast errors with varying lead times using ERA5 at 0 hour as inputs.
@@
Figure 8: Root mean square errors (RMSEs, shown by the numbers in the cell) of geopotential at 500hPa, temperature at 850hPa, and temperature at 2m from the 48-hour forecast using assimilated data as inputs, and from forecasts with lead times from 48-hour to 96-hour using ERA5 as inputs. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 96-hour forecast errors.
@@
As the scoreboard plot in Figure 8 shows, results from this experiment reveal that the 48-hour forecast errors using assimilated data inputs gradually converge to the 48-hour forecast errors using ERA5 inputs. More results can be found in subsection A.3. The errors using assimilated data cannot be lower than using ERA5 inputs because the ERA5 data is used to simulate observations in the experiment. However, this setup allows us to compare the forecast errors against ones with longer lead times to determine the extent of lead time lost due to the use of assimilated data as forecast inputs. With 8,000 observed columns, the errors for z500, t850, and t2m are lower than the 72-hour forecast errors. This implies that the lead time lost is less than 24 hours for these variables. With 40,000 observed columns, the error for z500 and t850 is approximately equivalent to the 60-hour forecast error. However, the t2m errors remain larger than the 60-hour forecast errors, potentially due to diurnal effects that cause a decrease in the forecast errors of t2m from the 48-hour lead time level to a 60-hour lead time level.
3.5 Limitations and future works
Currently, our method can consume point measurements at a single time step as observations (subsection 2.1). This covers weather station and upper-air balloon measurements but does not include satellite imagery and radar soundings which can be represented as functions of the atmospheric states. We are planning to support broader types of observations and multiple time steps using loss-function-based conditioning at the inference period [Chung et al., 2022, Rozet and Louppe, 2023a]. In addition, we are going to explore adding conditioning of the satellite imagery directly into the diffusion model similar to the conditioning of the predicted state (subsection 2.3). This requires much more training effort but provides more flexibility as the observation function is learned implicitly. With more input observations, we expect to have more accurate and stable results when performing autoregressive data assimilation.
On the other hand, our method lacks quality control of the input observation data. It takes the input observations as ground truth values and enforces assimilated data to be consistent with the observations. Therefore, we have to resort to simulated observations from the ERA5 dataset. We are going to investigate possible quality control techniques that either pre-process the data separately or be embedded in the data assimilation process. By addressing the above points, we will achieve a fully operational data assimilation model.
@@ 3.6 Conclusion
In summary, we propose a data assimilation method, DiffDA, capable of assimilating high-resolution atmospheric variables up to a horizontal resolution of ${}^{\circ}$, setting a new record for ML-based data assimilation models. Based on the denoising diffusion model, we adapt the pre-trained GraphCast neural network as the backbone model motivated by its compatible input and output shapes. DiffDA’s flexibility also allows for integrating other forecast models, ensuring easy updates and maintenance. A key feature of DiffDA is the enforcement of conditioning on the predicted state during training and inference periods, and the conditioning of sparse observations exclusively at the inference stage. An additional benefit of this conditioning approach is the automatic generation of a post-processing model, should observations not be supplied at inference time.
The experimental results validated the effectiveness of our method: the assimilated data converges to the observations as the number of observed columns increases. With less than $0.77\%$ of total “real-world” data columns and 48-hour forecast, the error of assimilated data is on par with 24-hour forecast error. When used as an input for forecast models, those assimilated data resulted in a maximum lead time loss of 24 hours compared with using the ERA5 dataset as inputs. This enables running data assimilation and simulation in an autoregressive cycle. It remains a challenge to constrain the errors across the autoregressive iterations. This is likely to be solved by assimilating more real-world observations.
All the data assimilation experiments can run on a single high-end PC with a GPU within 15 - 30 minutes per data assimilation step, while a similar task using traditional methods typically requires large compute clusters. This indicates a significant reduction in computational costs. It also opens up the possibility of assimilating more observational data that is otherwise discarded by traditional methods and producing assimilated data with higher accuracy.
4 Ethical and social impact
This work aims for a more efficient and accurate data assimilation method, which will potentially help reduce the energy costs of weather forecast centers, make better weather forecasts, and provide better data for tackling climate change.
Appendix A Appendices
A.1 Code Availability
The source code is available in this repository https://anonymous.4open.science/r/DiffDA-BED0. The repository includes a modified GraphCast model which has the Apache License version 2.0.
A.2 Scoreboard plots for single-step data assimilation
Figure 9: Root mean square errors (RMSEs, shown by the numbers in the cell) of surface variables from the single-step assimilated data, and from 6-hour to 48-hour GraphCast forecasts. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 48-hour forecast errors.
Figure 10: Root mean square errors (RMSEs, shown by the numbers in the cell) of geopotential from the single-step assimilated data, and from 6-hour to 48-hour GraphCast forecasts. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 48-hour forecast errors.
Figure 11: Root mean square errors (RMSEs, shown by the numbers in the cell) of temperature from the single-step assimilated data, and from 6-hour to 48-hour GraphCast forecasts. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 48-hour forecast errors.
Figure 12: Root mean square errors (RMSEs, shown by the numbers in the cell) of specific humidity from the single-step assimilated data, and from 6-hour to 48-hour GraphCast forecasts. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 48-hour forecast errors.
Figure 13: Root mean square errors (RMSEs, shown by the numbers in the cell) of horizontal wind speed (U) from the single-step assimilated data, and from 6-hour to 48-hour GraphCast forecasts. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 48-hour forecast errors.
Figure 14: Root mean square errors (RMSEs, shown by the numbers in the cell) of horizontal wind speed (V) from the single-step assimilated data, and from 6-hour to 48-hour GraphCast forecasts. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 48-hour forecast errors.
Figure 15: Root mean square errors (RMSEs, shown by the numbers in the cell) of vertical wind speed from the single-step assimilated data, and from 6-hour to 48-hour GraphCast forecasts. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 48-hour forecast errors.
A.3 Scoreboard plots for forecast on single-step assimilated data
Figure 16: Root mean square errors (RMSEs, shown by the numbers in the cell) of surface variables from the 48-hour forecast using assimilated data as inputs, and from forecasts with lead times from 48-hour to 96-hour using ERA5 as inputs. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 96-hour forecast errors.
Figure 17: Root mean square errors (RMSEs, shown by the numbers in the cell) of geopotential from the 48-hour forecast using assimilated data as inputs, and from forecasts with lead times from 48-hour to 96-hour using ERA5 as inputs. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 96-hour forecast errors.
Figure 18: Root mean square errors (RMSEs, shown by the numbers in the cell) of temperature from the 48-hour forecast using assimilated data as inputs, and from forecasts with lead times from 48-hour to 96-hour using ERA5 as inputs. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 96-hour forecast errors.
Figure 19: Root mean square errors (RMSEs, shown by the numbers in the cell) of specific humidity from the 48-hour forecast using assimilated data as inputs, and from forecasts with lead times from 48-hour to 96-hour using ERA5 as inputs. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 96-hour forecast errors.
Figure 20: Root mean square errors (RMSEs, shown by the numbers in the cell) of horizontal wind speed (U) from the 48-hour forecast using assimilated data as inputs, and from forecasts with lead times from 48-hour to 96-hour using ERA5 as inputs. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 96-hour forecast errors.
Figure 21: Root mean square errors (RMSEs, shown by the numbers in the cell) of horizontal wind speed (V) from the 48-hour forecast using assimilated data as inputs, and from forecasts with lead times from 48-hour to 96-hour using ERA5 as inputs. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 96-hour forecast errors.
Figure 22: Root mean square errors (RMSEs, shown by the numbers in the cell) of vertical wind speed from the 48-hour forecast using assimilated data as inputs, and from forecasts with lead times from 48-hour to 96-hour using ERA5 as inputs. The errors are calculated against the ERA5 data. The cells are color-coded with the RMSEs relative to the 96-hour forecast errors.